OpenStudy (anonymous):

In a bag containing black and white marbles,3/8 of the marbles are black. Which could be the numbers of black and white marbles in the bag? A. 3 black and 8 white B. 12 black and 20 white C. 20 black and 12 white D. 8 black and 3 white @Pawtpie

OpenStudy (anonymous):

So 3/8 is the ratio of black marbles to total marbles

OpenStudy (elonasushchik):

Option B. total = 32 black 12 so it is 8/3

OpenStudy (anonymous):

can u please work it out so I understand it

OpenStudy (elonasushchik):

Medal and fan?! Thx

OpenStudy (anonymous):

I don't think B is correct

OpenStudy (anonymous):

B. there are 3 black and 5 white. 3*4=12 and 5*4=20.

OpenStudy (anonymous):

Since 3/8 is the ratio of black marbles to the total, 3/5 is the ratio of black marbles to right

OpenStudy (anonymous):

white*

OpenStudy (elonasushchik):

No@Pawtpie it's B. I just did this and it was b

OpenStudy (anonymous):

I'm sorry B looks correct

OpenStudy (mathstudent55):

Each answer tells you a number of black marbles and a number of white marbles. For each answer, add the number of black marbles and white marbles. Then write a ratio of the number of black marbles to the total and compare with 3/8.

OpenStudy (elonasushchik):

Throw me the medal?! Thx

OpenStudy (anonymous):

yep i think its B too. if you divide both of the numbers 12 and 20 by 4 (the number of times they drew) it equals 3 and white and 5 black

OpenStudy (anonymous):

*3 white

OpenStudy (mathstudent55):

Example: Choice A. 3 black & 8 white Total is 3 + 8 = 11 Ratio of black to total is: 3/11 which is not 3/8, so A is not the answer. Now do the same for the other choices.

OpenStudy (anonymous):

@Pawtpie I got 100% thank you all

OpenStudy (elonasushchik):

Your welcome Shelby

OpenStudy (anonymous):

u got the medal

OpenStudy (elonasushchik):

Thx